OpenStudy (anonymous):

algebra...HELPPP 6.4x + 2 = 3.2 - 4x Im trying to find what x is. could you show me the method of how you got the answer aswell. Myne keeps coming up different to the answer :( Thanks for the help also :)

OpenStudy (crashonce):

move x to one side and number sto other

OpenStudy (crashonce):

6.4x + 4x = 3.2-2 therefore 10.4x = 1.2 104x = 12 26x = 3 x = 3/26

OpenStudy (anonymous):

why did you take it out of decimal form?

OpenStudy (anonymous):

@CrashOnce

OpenStudy (crashonce):

times by 10 on both sides, makes the thing easier

OpenStudy (anonymous):

thank you so much :)